Find the solution of the given initial value problem. (4) – 6y + 9 = 0 y(1) = 10 + e, y (1) = 8 + 3e3, y (1) = 9e), y
Answer #1

The value = 0 The equation 2 son The value Then, 3 given ini tial problem is yu) - 6y + 9 y - auxiliary 6m3 + 9 m²= m² 6m +98 we get We get 3 Solving equation and (4) we get 3 C3 + 5 C4 3 C4+ C) = 3-3 = 0 C4 = = CH = 0 Then C3=1. Putting C3=1 and CVerification: dorg ъс Here, ч (1) = 2+ 8 x + Then a(х) - 3 3 Now, Ч (1) 9e3x A). So, 5 27 е lot and 9 ( x) 4) ) = 91 с 2 N